please help me solve by the substitution method this problem x + y = 6
y = x + 2 Thank you

Answer by Alan3354(69443)

x + y = 6
y = x + 2
---------
Sub for y into the 1st eqn
x + (x+2) = 6
2x + 2 = 6
2x = 4
x = 2
---------
y = x+2
y = 4